calculate the rms velocity chlorine at 17 C and 800 mm pressure
 

Grade:11

1 Answers

Arun
4 years ago
Root-mean-square. speed of an ideal gas: 
v (rms) = sqrt(3RT/M) 
where R = 8.31 J K−1·mol−1 is the gas constant, T is the absolute temperature, M the molar mass. 
T = 17 + 273 = 290°K 
M(chlorine) = 0.071 kg·mol−1 

v (rms) = sqrt(3×8.31×290/0.071) = 306.53m/s